OpenStudy (anonymous):

sonia read 2 pages of a book per minute. If Sonia read 2 pages of the book at the same rate, which equation shows the total number of pages,y, that sonia read in x minute Please help! I have school tomorrow! if you can i will fan and medal you!

OpenStudy (anonymous):

19. Sonia read 2 pages of a book per minute. If Sonia read all the pages of the book at the same rate, which equation shows the total number of pages, y, that Sonia read in x minutes? 20. A x = 2y B y = 2x C x =2+y D y =2+x

OpenStudy (anonymous):

Sorry copy and paisted wrong thing! @jabberjaw thats it ^

OpenStudy (anonymous):

Inequalities are not my strong point. I am NOT positive if this is the answer, but I would say B.

OpenStudy (anonymous):

We want to know the total number of pages: y = 2(x) Sonia reads 2 pages (per) minute. Key word is per. That usually means multiplication.

OpenStudy (anonymous):

ok! thank you sooooo much that really helped! @jabberjaw